OpenStudy (anonymous):

1 MEDAL AND FAN QUESTION Look at the following number sequence. 3, 11, 19, 27, . . . Based on the pattern, what are the next two terms? 31, 40 35, 43 216, 1728 43, 51

OpenStudy (anonymous):

3 + 8 = 11 + 8 = 19 + 8 = 27 + 8 = 35 + 8 = 35 B) 35, 43

OpenStudy (anonymous):

THANK YOU

OpenStudy (anonymous):

You're Very Welcome